11) On a clear day at a certain location, a 96 V / m vertical electric field exists near theEarth's surface. At the same place, the Earth's magnetic field has a magnitude of 6.20 X10-5 T. Compute the energy densities of the two fields.
uE = ___________ nJ/m3
uB= ___________ µJ/m3
Answer #1
the energy density :
electric field :
U(E) = (1/2)εo E^2 = (1/2)*8.85x10^-12*96^2 = 40.78 nJ/m3
magnetic field :
U(B) = (1/2)*B^2/μo = (1/2)*(6.2x10^-50^2/(4πx10^-7) = 1529.5 μJ/m3